1.a. Imagine that you are flipping 3 coins. How many possible outcomes (arrangement of coins) can there be when flipping 3 coins? (Think outcomes not microstates)

_____ outcomes
Answer #1

answer : 8 outcomes

Explanation :

there are 3 coins flipping ,

each coin have 2 options , 1) head , 2) tails

so

number of outcomes = 2^n

n = number of coins

outcomes = 2^3 = 8

possible outcomes = 8